OpenStudy (anonymous):

Show all work & use complete sentences to explain. : 4t^2-1/15t+10divided by 4t^2-4t+1/3t^2+5t+2

OpenStudy (anonymous):

@arjont Help please?

OpenStudy (anonymous):

This is similar to the previous question, again make sure to factor out each term, and then try to see if you can simplify the expression. make sure to note the restrictions before you make any simplifications. remember that dividing two fractions is the same as multiplying 1st one by the reciprocal of the 2nd one, that is \[\frac{ a/b }{ c/d }=\frac{a}{b}\frac{d}{c}\] \[\frac{ 4t^2-1 }{15t+10 }\frac{3t^2+5t+2}{4t^2-4t+1} = \frac{(2t-1)(2t+1)(t+1)(3t+2)}{5(3t+2)(2t-1)^2}\] you may simplify the expression further if you wish (or if it asks you to do so)